At the begining of a walk, Roberto and Juaita are 6.6 miles apart. If they leave at the same time and wal in the same direction, Roberto overtakes Juanita in 11hours. If they wal towards each other, they meet in 1 hour. What are their speeds?
Answer by mananth(16949)

Roberto speed = xmph
Juanita speed = y mph
same direction so subtract speed
effective speed = x-y
time = d/r
6.6/(x-y)=11
11x-11y=6.6.....1

Opposite direction
add up speeds
effective speed = x+y
time = 6.6/(x+y)
6.6/(x+y)=1
x+y=6.6..........2
11x-11y =6.6 .............1
x+y=6.6 .............2

multiply (1)by 1
Multiply (2) by 11
11x-11y=6.6
11x+11y =72.6
Add the two equations
22x=79.2
/22
x=3.60 mph
plug value of x in (1)
11x-11y=6.6
39.6-11y=6.6
-11y=6.6-39.6
-11y=-33
y=3.00 mph